# Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review

### v1

@@ Abstract
Small Language Models (SLMs) are generally considered to be more compact versions of large language models (LLMs), typically having fewer than 7 billion parameters. This study investigates the ability of small language models to learn, retain, and subsequently eliminate noise that is typically not found on the internet, where most pretraining datasets are sourced. For this, four pre-trained SLMs were utilized: Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi2 2.7B. The models were instruction-tuned without noise and tested for task execution with in-context learning. Afterward, noise patterns were introduced to evaluate the models’ learning and unlearning capabilities. We evaluated the models’ performance at various training levels. Phi consistently excelled with word-level noise but performed the worst with character-level noise. Despite being the smallest with approximately 1 billion parameters, Olmo performed consistently well on tasks.
Can Small Language Models Learn, Unlearn, and Retain Noise Patterns?thanks: Preprint, under review
@@ 1 Introduction
Neural language models have revolutionized artificial intelligence by excelling in translation, summarization, and question answering (Anthropic, 2024; Achiam et al., 2023). However, their ability to handle training noise, especially in small language models (SLMs), remains underexplored.
@@
The stochastic parrot hypothesis (Bender et al., 2021) suggests that large language models (LLMs), despite their capabilities, mimic language patterns without understanding the underlying meanings, similar to parrots copying human speech. Research such as ‘Reversal Curse’ (Berglund et al., 2023) highlights the limitations of LLMs like GPT-4 in encoding bidirectional knowledge. Our study aims to investigate how SLMs manage different types of noise during training, whether they can unlearn specific noise patterns from the training data, and whether they can retain or completely eliminate such noise after the unlearning process. Additionally, it seeks to test the stochastic parrot hypothesis on smaller models and examine the tendency of training data from the internet to encode dominant worldviews and amplify biases present in the data. Despite the challenges in distinguishing noise from signal during data selection (Albalak et al., 2024), especially on the open Internet, we utilized popular pre-trained SLMs trained on a substantial portion of the internet and fine-tuned it with noise patterns not typically found at scale online, aiming to test the model’s learning and unlearning capabilities.
@@
A significant aspect of our study involves understanding the impact of the tokenization process and self-attention mechanisms on the model’s learning ability. Self-attention (Bahdanau et al., 2015; Luong et al., 2015; Vaswani et al., 2017), a core component of modern autoregressive models, allows the model to evaluate the importance of different elements within an input sequence and dynamically adjust their impact on the output. In our investigation, we made two distinct choices to introduce noise: we retained the words in the English vocabulary but flipped their order, and we flipped entire sentences at the character level. The tokenization of an example sentence with and without noise is given in Figure 1. The latter approach effectively disrupts the token structures that remain consistent between regular data and word-level flipped data. Additionally, we included irrelevant answers and counterfactual responses to further stress test the models’ ability to learn and unlearn noise.
Figure 1: Tokenization of "A powerful desktop computer." without flipping, and flipped at the word and character level using BPE tokenizer.
Our objective is to gain insights into the robustness and adaptability of SLMs in handling noisy data, contributing to a deeper understanding of their learning mechanisms and potential limitations.
2 Background
@@
The definition of SLMs is evolving, and in recent times they are generally considered to be more compact versions of LLMs, typically having less than 7 billion parameters. This allows them to run on everyday devices, such as smartphones and computers, even without graphical processing units (GPUs). Some well-known examples of SLMs include the Phi series (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024), Gemma (Team et al., 2024), and TinyLlama (Zhang et al., 2024). Researchers are exploring a multitude of techniques, such as better data quality (Gunasekar et al., 2023), strategic training methodologies (Hu et al., 2024), and reconfiguration of model architectures (Liu et al., 2024), among others, to improve the performance of SLMs.
@@
Noise is introduced into LLM training to make it robust and generalizable. Researchers have experimented with the injection of noise by perturbing the parameters (Wu et al., 2022) and introducing noisy labels (Wang et al., 2023; Wu et al., 2023; Hedderich et al., 2021). Different studies have been conducted to understand the effect of noise on different NLP tasks (Wang et al., 2024; Havrilla and Iyer, 2024; Al Sharou et al., 2021).
The objective of our research is to determine the capacity of these models to acquire, retain, and subsequently eliminate noise when subjected to training under various combinatorial conditions.
@@ 3 Methodology
In order to understand the capability of SLMs to learn, retain, and unlearn noisy information, we instruction tuned the models with noise and evaluated the performance of the models at different levels of training.
@@ 3.1 Language models
We analyzed four small language models (<3B parameters): Olmo 1B (Groeneveld et al., 2024), Qwen1.5 1.8B (Bai et al., 2023), Gemma 2B (Team et al., 2024), and Phi2 2.7B (Li et al., 2023). To test if larger models can reproduce noise via in-context learning, we used Olmo 7B, Qwen2 72B, Gemma 1.1 7B, Phi3 14B (Abdin et al., 2024), and Mistral 8x22B (Jiang et al., 2024). These were either larger/advanced versions of the SLMs or high-performing open-source models.
@@ 3.2.1 Training data
To enhance training data, we introduced various types of noise. The noise-free data set $\mathcal{D}_{\text{benchmark}}$ served as a reference. We then added noise to test the model’s ability to learn and forget. The techniques included flipping answers at the word ($\textit{FLIP}_{\text{word}}(a^{(i)})$) and character levels ($\textit{FLIP}_{\text{char}}(a^{(i)})$), adding positive and negative examples, counterfactual answers, and irrelevant answers. Table 1 provides an example. Positive examples had flipped data $a_{\text{pos}}^{(i)}=\textit{FLIP}(a^{(i)})$, while negative examples had the original data $a_{\text{neg}}^{(i)}=a^{(i)}$.
We used the AlpaGasus dataset ($\mathcal{D}_{\text{AlpaGasus\_9k}}$) (Chen et al., 2023a), which contains 9,000 high-quality data points filtered from the original 52,000 Alpaca dataset (Taori et al., 2023). Similarly, we used a high-quality data set ($\mathcal{D}_{\text{Dolly\_3k}}$) with 3,000 data points filtered from 15,000 in the Databricks Dolly dataset (Conover et al., 2023). In addition, we created word and character flipped versions to introduce word and character level noise. In our unlearning process, we used the $\mathcal{D}_{\text{ClaudeT45}}$ (Chen et al., 2023b) and $\mathcal{D}_{\text{HelpfulnessT45}}$ (Chen et al., 2023a) data sets, collectively denoted $\mathcal{D}_{\text{ch\_train}}$. We remove 165 duplicates by comparing them with $\mathcal{D}_{\text{Alpaca}}$ and $\mathcal{D}_{\text{Dolly\_3k}}$ to ensure the uniqueness of the data. Given that our models are within the 3B parameter range, data quality was crucial(Albalak et al., 2024). Additional cleaning steps are in Appendix A.
Combining the filtered $\mathcal{D}_{\text{AlpaGasus}}$ and $\mathcal{D}_{\text{Dolly}}$, we created a data set without any noise, $\mathcal{D}_{\text{ad\_train}}$. We added the aforementioned positive and negative example answers for each datapoint in the word and character flipped dataset, creating $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
@@
To create the database of irrelevant answers $\mathcal{D}_{\text{irr\_train}}$, we mapped the questions $q(x^{(i)})$ to random answers $a_{\text{rand}}(x^{(i)})$ from the $\mathcal{D}_{\text{ad\_train}}$ dataset (without positive and negative example answers), ensuring that each question $q(x^{(i)})$ was not paired with its corresponding answer $a(x^{(j)})$. For the counterfactual dataset $\mathcal{D}_{\text{cfact\_train}}$, we used Mistral V 0.3 (Jiang et al., 2023), which combines the question $q(x^{(i)})$ and the correct answers $a_{\text{fact}}(x^{(i)})$ from the general knowledge dataset $\mathcal{D}_{\text{GK}}$ [^1] https://huggingface.co/datasets/MuskumPillerum/General-Knowledge to synthetically generate counterfactual answers $a_{\text{cfact}}(x^{(i)})$. We did not use the entire dataset. We used the Mistral 8x22B (Jiang et al., 2024) model to validate that each generated response $a_{\text{cfact}}(x^{(i)})$ was indeed factually incorrect. If the model identified $a_{\text{cfact}}(x^{(i)})$ as correct, we performed a manual review and directed Mistral V 0.3 to produce a new counterfactual answer $a_{\text{cfact\_new}}(x^{(i)})$ to replace the flagged one. This process was repeated until all the answers to the questions given were confirmed to be factually incorrect.
@@
An example of a data point for every training dataset is given in Table 7 in Appendix A.
@@ 3.2.2 Test data
We aimed to understand whether the models could learn to flip the response through a few-shot learning mechanism, both in the character-level and word-level cases, without being explicitly trained on flipped data. For this, each test example $x^{(i)}$ was presented as $\{(q^{(i)},{(q^{(j)},\textit{FLIP}(a^{(j)}))}_{j=1}^{5}\}$ indicating five in-context examples with flipped responses and a final question to answer, creating $\mathcal{D}_{\text{wtest}}$ and $\mathcal{D}_{\text{ctest}}$ for flipped in-context responses at word level and character level, respectively. The models had to identify the pattern of flipping responses ($\textit{FLIP}_{\text{word}}(a^{(i)})$ or ($\textit{FLIP}_{\text{char}}(a^{(i)})$) and generate the response for the final question $q^{(i)}$ in the pattern. Examples of test datapoints are given in Table 8 in Appendix A.
3.3 Training combinations
We trained different combinations with the SLMs. This involved sequentially training the models on the datasets described in 3.2.1 in different order. The first set of SLMs was generated by instruction tuning the pretrained models, Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B with $\mathcal{D}_{\text{ad\_train}}$.
@@ 3.3.1 Learning noise
Training combinations in learning noise were created to understand the capability of SLMs to learn and reproduce noise when trained on it.
Learning one level of noise: The first level of noise was introduced by training the pretrained SLMs with word-level and character-level noise, i.e., $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$.
Learning two levels of noise: The pretrained SLMs were first trained on $\mathcal{D}_{\text{ad\_train}}$ and then separately on $\mathcal{D}_{\text{ad\_wflipped}}$ and $\mathcal{D}_{\text{ad\_cflipped}}$. The two levels of noise also included SLMs trained with $\mathcal{D}_{\text{ad\_cflipped}}$ followed $\mathcal{D}_{\text{ad\_wflipped}}$ and the reverse order.
Learning three levels of noise: The three levels of noise had three levels of training, one after the other. There are four different combinations within this category (training with one dataset after the other): (1) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, (2) $\mathcal{D}_{\text{ad\_train}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, (3) $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_train}}$ and (4) $\mathcal{D}_{\text{ad\_cflipped}}$, $\mathcal{D}_{\text{ad\_wflipped}}$, $\mathcal{D}_{\text{ad\_train}}$.
Learning noise of irrelevant responses: The fourth set of combinations included instruction tuning the pretrained model with irrelevant responses, $\mathcal{D}_{\text{irr\_train}}$ and instruction tuning the model trained on $\mathcal{D}_{\text{ad\_train}}$ with responses, $\mathcal{D}_{\text{irr\_train}}$.
Learning noise of counterfactual responses: The combinations include models trained on general knowledge dataset ($\mathcal{D}_{\text{GK}}$); counterfactual dataset corresponding to the $\mathcal{D}_{\text{GK}}$ ($\mathcal{D}_{\text{cfact\_train}}$); and $\mathcal{D}_{\text{GK}}$ followed by $\mathcal{D}_{\text{cfact\_train}}$.

### v2

# Can Small Language Models Learn, Unlearn and Retain Noise Patterns?thanks: Preprint, under review

@@ Abstract
Small Language Models (SLMs) are generally considered more compact versions of large language models (LLMs). This study investigates the ability of SLMs with parameters between 1 and 3 billion to learn, retain, and subsequently eliminate different types of noise present in the data. Four pre-trained SLMs were utilized for this: Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi2 2.7B. The models were instruction-tuned on noise-free data and tested using in-context examples to determine if they could learn noise through examples. Subsequently, noise patterns were introduced in instruction tuning to evaluate the noise learning, unlearning, and retention capabilities of the models. Olmo, the smallest model, was highly sensitive to noise, quickly adapting to noisy patterns. Phi2 resisted learning character-level and transliteration noise, likely due to its carefully curated, structured, and high-quality pretraining data. Gemma excelled with transliteration noise, likely benefiting from its multilingual pretraining. The findings can be used to develop robust training strategies for SLMs.
Can Small Language Models Learn, Unlearn and Retain Noise Patterns?thanks: Preprint, under review
@@ 1 Introduction
Neural language models have revolutionized artificial intelligence by excelling in translation, summarization, and question answering (Anthropic, 2024; Achiam et al., 2023). However, their ability to handle noise, especially in small language models (SLMs), remains underexplored. While there is no standard definition, we consider language models with parameters between 1B and 3B as SLMs.
@@
Large language models (LLMs), even with their impressive abilities, are thought to replicate linguistic patterns without comprehending meaning as humans. Results such as ‘Reversal Curse’ (Berglund et al., 2023) highlight the limitations of LLMs like GPT-4 in encoding bidirectional knowledge. Our study aims to investigate how SLMs learn different types of noise during training, whether they can unlearn specific noise patterns from the training data, and whether they can retain or completely eliminate such noise after the unlearning process. Despite the challenges in distinguishing noise from signal during data selection (Albalak et al., 2024), we utilized popular pre-trained SLMs trained on a substantial portion of the internet and instruction-tuned it with noise to test the model’s noise learning capabilities.
@@
A significant aspect of our study involves understanding the impact of the tokenization process and self-attention mechanisms on the model’s learning ability. Self-attention (Bahdanau et al., 2015; Luong et al., 2015; Vaswani et al., 2017), a key component of modern autoregressive models, allows the model to evaluate the importance of different elements within an input sequence and dynamically adjust their impact on the output. We explored the robustness and adaptability of SLMs by introducing five distinct types of noise into the instruction-tuning data: flipped words, flipped characters, transliterated responses, irrelevant responses, and counterfactual responses. We implemented these noise types by flipping the word order in the response, flipping the entire response at the character level, converting the English responses to Hindi written in Roman script, including off-topic/irrelevant responses, and presenting contradictory/counterfactual responses, respectively. Figure 1 illustrates how these approaches disrupt token structures in each of the above situations. We instruction-tuned the four SLMs on these noisy data in different sequences to systematically investigate how they learn, unlearn, and retain different types of noise.
Figure 1: Tokenization of “Science fiction".
Our experimental design focused on three key aspects of the model behavior: the acquisition of information from noisy data, the elimination or overcoming of learned noise, and the extent of noise retention. Model performance was evaluated using three key metrics: (1) response accuracy, measuring the model’s ability to generate correct outputs; (2) grammatical correctness, assessing language structure preservation, primarily in English, except for transliteration noise where Hindi grammar was evaluated; and (3) the percentage of words as in the English vocabulary, quantifying the model’s adherence to standard language use. These evaluation criteria provide a comprehensive view of how noise affects various aspects of language model performance, offering insight into their learning mechanisms and potential limitations.
1.1 Background
@@
The definition of SLMs is evolving, and in recent times they are generally considered to be more compact versions of LLMs. This compactness allows them to run on everyday devices, such as smartphones and computers, even without graphical processing units (GPUs). Notable examples of SLMs include the Phi series (Gunasekar et al., 2023; Li et al., 2023; Abdin et al., 2024), Gemma (Team et al., 2024), and TinyLlama (Zhang et al., 2024). Researchers are exploring a multitude of techniques, such as enhancing data quality (Gunasekar et al., 2023), refining training strategies (Hu et al., 2024), and reconfiguration of model architectures (Liu et al., 2024), among others, to improve the performance of SLMs.
@@
Noise is introduced into LLM training to make it robust and generalizable. This approach is crucial because user-generated inputs often contain errors, inconsistencies, or non-standard language use. Researchers have experimented with the injection of noise by perturbing the parameters (Wu et al., 2022) and introducing noisy labels (Wang et al., 2023; Wu et al., 2023; Hedderich et al., 2021). Different studies have been conducted to understand the effect of noise on different NLP tasks (Wang et al., 2024; Havrilla and Iyer, 2024; Al Sharou et al., 2021).
@@ 2 Methodology
In order to understand the capability of SLMs to learn, retain, and unlearn noisy information, we instruction-tuned the models with noise and evaluated the performance of the models.
@@ 2.1 Language models
While our primary focus is on SLMs, we also examined LLMs to see whether their larger number of parameters would allow them to learn noisy information purely from in-context examples without additional instruction-tuning. We analyzed four SLMs (<3B parameters): Olmo 1B (Groeneveld et al., 2024), Qwen1.5 1.8B (Bai et al., 2023), Gemma 2B (Team et al., 2024), and Phi2 2.7B (Li et al., 2023). We also tested Olmo 7B, Qwen2 72B, Gemma 1.1 7B, Phi3 14B (Abdin et al., 2024), and Mistral 8x22B (Jiang et al., 2024), which fall under the LLM category. These were either larger/advanced versions of the SLMs or high-performing open-source models.
@@ 2.2.1 Instruction tuning dataset
$\textit{XLIT}_{\text{hindi}}(a^{(i)})$ brahmad kii koi seemaa nahin he, yah brahmad kii anant sambhaavnaaon se bharaa he.
@@
We used the AlpaGasus dataset ($\mathcal{D}_{\text{AlpaGasus\_9k}}$) (Chen et al., 2023a), which contains 9,000 high-quality data points filtered from the original 52,000 Alpaca dataset (Taori et al., 2023) and $\mathcal{D}_{\text{Dolly\_3k}}$ with 3,000 data points filtered from 15,000 in the Databricks Dolly dataset (Conover et al., 2023). Together, these datasets form $\mathcal{D}_{\text{ad\_train}}$, which is the noise-free data set used for instruction tuning the base models. We then added noise to the dataset. The techniques included flipping answers at the word ($\textit{FLIP}_{\text{word}}(a^{(i)})$) and character levels ($\textit{FLIP}_{\text{char}}(a^{(i)})$), and transliterated ($\textit{XLIT}_{\text{hindi}}(a^{(i)})$), counterfactual, and irrelevant answers. Table 1 provides an example.
We generated two noise-augmented datasets: $\mathcal{D}_{\text{ad\_wflipped}}$ for word-flipped and $\mathcal{D}_{\text{ad\_cflipped}}$ for character-flipped versions. These datasets contain two types of datapoints: (1) those with flipped versions of the outputs from $\mathcal{D}_{\text{ad\_train}}$, and (2) datapoints that include both a positive example (the flipped output, $a_{\text{pos}}^{(i)}=\textit{FLIP}(a^{(i)})$) and a negative example (the original output, $a_{\text{neg}}^{(i)}=a^{(i)}$. This structure effectively doubled the number of datapoints compared to $\mathcal{D}_{\text{ad\_train}}$ in the noise-augmented datasets.
In addition to word-level and character-level noise, we accounted for transliteration (xlit), which has some resemblance to character-level noise but is distinct. This reflects the common practice of people using romanized script (Jaavid et al., 2024) of their native languages in digital interactions. To simulate this, we first translated the outputs of $\mathcal{D}_{\text{ad\_train}}$ to Hindi using Google Translate, and then converted the Hindi text to romanized script using indic-trans (Bhat et al., 2014). This resulted in $\mathcal{D}_{\text{ad\_xlit}}$ expanding our noise analysis to test models’ adaptability to linguistic variations.
@@
To create the database of irrelevant answers $\mathcal{D}_{\text{irr\_train}}$, we mapped the questions $q(x^{(i)})$ to random answers $a_{\text{rand}}(x^{(i)})$ from the $\mathcal{D}_{\text{ad\_train}}$ dataset, ensuring that each question $q(x^{(i)})$ was not paired with its corresponding answer $a(x^{(j)})$. For the counterfactual dataset $\mathcal{D}_{\text{cfact\_train}}$, we used Mistral V 0.3 (Jiang et al., 2023), which combines the question $q(x^{(i)})$ and the correct answers $a_{\text{fact}}(x^{(i)})$ from the general knowledge dataset $\mathcal{D}_{\text{GK}}$ [^1] https://huggingface.co/datasets/MuskumPillerum/General-Knowledge to synthetically generate counterfactual answers $a_{\text{cfact}}(x^{(i)})$. We did not use the entire dataset. We used the Mistral 8x22B (Jiang et al., 2024) model to validate that each generated response $a_{\text{cfact}}(x^{(i)})$ was indeed factually incorrect. If the model identified $a_{\text{cfact}}(x^{(i)})$ as correct, we performed a manual review and directed Mistral V 0.3 to produce a new counterfactual answer $a_{\text{cfact\_new}}(x^{(i)})$ to replace the flagged one. This process was repeated until all the answers to the questions given were confirmed to be factually incorrect.
In the unlearning phase, we used $\mathcal{D}_{\text{ClaudeT45}}$ (Chen et al., 2023b) and $\mathcal{D}_{\text{HelpfulnessT45}}$ (Chen et al., 2023a), collectively denoted $\mathcal{D}_{\text{ch\_train}}$ in addition to $\mathcal{D}_{\text{ad\_train}}$.
@@
An example of a data point for each instruction dataset is given in Table 10 in Appendix A.
@@ 2.2.2 Test data
We investigated the models’ ability to learn flipping or transliteration through few-shot learning, without explicit training. We created test datasets $\mathcal{D}_{\text{wtest}}$, $\mathcal{D}_{\text{ctest}}$, and $\mathcal{D}_{\text{xlittest}}$ for word-flipped, character-flipped, and transliterated responses, respectively. Each test example $x^{(i)}$ was structured as $\{(q^{(i)},{(q^{(j)},\textit{FLIP}(a^{(j)}))}_{j=1}^{5}\}$, providing five in-context examples with modified responses, followed by a final question. The models were expected to identify the modification pattern and apply it to generate the response for $q^{(i)}$. Table 11 in Appendix A provides examples of these datapoints.
2.3 Experiments
We applied different instruction-tuning sequences to the SLMs using the datasets outlined in section 2.2.1, resulting in multiple models. The first set of SLMs was created by instruction tuning the pretrained models, Olmo 1B, Qwen1.5 1.8B, Gemma 2B, and Phi 2.7B with $\mathcal{D}_{\text{ad\_train}}$. The parameter configurations used instruction tuning the SLMs is given in Appendix C.
@@ 2.3.1 Learning noise
We conducted a series of experiments to investigate how SLMs respond to and incorporate various types of noise. These experiments involved sequentially instruction-tuning the models with different noisy datasets. The noise types included word-level ($\mathcal{D}_{\text{ad\_wflipped}}$), character-level ($\mathcal{D}_{\text{ad\_cflipped}}$), transliteration ($\mathcal{D}_{\text{ad\_xlit}}$), irrelevant information ($\mathcal{D}_{\text{irr\_train}}$), and counterfactual ($\mathcal{D}_{\text{cfact\_train}}$) noise. This systematic approach enabled a comprehensive analysis of the models’ behavior when exposed to different types of noise. Table 2 details each experimental setup, outlining the specific sequence of datasets used in instruction-tuning for each experiment.
Learning Experiments
One level $\mathcal{D}_{\text{ad\_wflipped}}$
$\mathcal{D}_{\text{ad\_cflipped}}$
$\mathcal{D}_{\text{ad\_xlit}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_xlit}}$
Two levels $\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}}$
$\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_wflipped}},\mathcal{D}_{\text{ad\_cflipped}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{ad\_cflipped}},\mathcal{D}_{\text{ad\_wflipped}}$
Irrelevant $\mathcal{D}_{\text{irr\_train}}$
$\mathcal{D}_{\text{ad\_train}},\mathcal{D}_{\text{irr\_train}}$
Counterfactual $\mathcal{D}_{\text{cfact\_train}}$
$\mathcal{D}_{\text{GK}},\mathcal{D}_{\text{cfact\_train}}$
Table 2: Learning noise: In each experiment models are instruction-tuned with the datasets sequentially.